OpenStudy (anonymous):

The sum of the digits of a two-digit counting number is 13. When the digits are reversed, the new number is 45 greater than the original number. What was the original number? Please show work.

OpenStudy (anonymous):

x + y = 13 10x + y = 10y + x + 45 x = 13-y so 10(13-y) + y = 10y + 13 -y + 45 130 - 13 - 45 = 10y -y - y 72 = 8y y = 9 so x = 13-9 = 4 the original number was 49